### v1

@@ Abstract
Driven by recent advances in vision language models (VLMs) and egocentric perception research, we introduce the concept of an egocentric procedural AI assistant (EgoProceAssist) tailored to step-by-step support daily procedural tasks in a first-person view. In this work, we start by identifying three core tasks: egocentric procedural error detection, egocentric procedural learning, and egocentric procedural question answering. These tasks define the essential functions of EgoProceAssist within a new taxonomy. Specifically, our work encompasses a comprehensive review of current techniques, relevant datasets, and evaluation metrics across these three core areas. To clarify the gap between the proposed EgoProceAssist and existing VLM-based AI assistants, we introduce novel experiments and provide a comprehensive evaluation of representative VLM-based methods. Based on these findings and our technical analysis, we discuss the challenges ahead and suggest future research directions. Furthermore, an exhaustive list of this study is publicly available in an active repository that continuously collects the latest work: https://github.com/z1oong/Building-Egocentric-Procedural-AI-Assistant
@@
Egocentric perception, procedural tasks, AI assistant, VLMs.
@@ 1 Introduction
Procedural tasks, which require adherence to specific sequences of steps, are central to numerous daily and industrial activities. Errors in the order or execution of these steps can lead to inefficiencies, suboptimal outcomes, or even safety risks. Traditional video analysis methods, while capable of visual pattern recognition, lack a deep understanding of video content. They struggle to comprehend logical procedural narratives and face challenges in integrating multifunctional assistance, such as detection and video question answering (VQA). Recent advances in VLM [cheng2024videollama][lin2023video][li2024llava] and egocentric (first-person view) vision technologies [thatipelli2025egocentric][li2025challenges], combined with augmented and virtual reality devices, have enabled the development of egocentric AI assistants for procedural task guidance [yang2025egolife, huang2024vinci].
@@
Based on the envisioned functions, such as the ability to detect procedural errors in real time, the ability to learn the key steps of procedural tasks independently from videos, the ability to understand egocentric videos and answer relevant questions, this paper identifies three new technical approaches, to address key requirements for building an egocentric procedural AI assistant (EgoProceAssist): egocentric procedural error detection, egocentric procedural learning, egocentric procedural question answering. The overall structure is shown in Fig. 2.
Figure 1: Taxonomy of EgoProceAssist [flaborea2024prego, plini2024ti, huang2025modeling, patsch2025technical, lee2024error][ding2023every][storks2024transparent, haneji2024egooops, mazzamuto2025gazing, bansal2024united, alayrac2016unsupervised, elhamifar2019unsupervised, kukleva2019unsupervised, chowdhury2024opel, sener2015unsupervised, lin2022learning, zhou2018towards, zhukov2019cross, naing2020procedure, bansal2022my, mahmood2025procedure, shah2023steps, wang2023lifelongmemory, zhang2024hcqa, zhang2025hcqa, taluzzi2025pixels][yang2025egolife][di2024grounded][huang2024vinci][suglia2024alanavlm, chen2025grounded, biswas2025raven, ye2024mm]
The main contributions of this work can be summarized as follows:
@@
$\bullet$ We present novel taxonomies for methods in the three domains as shown in Fig. 1, enabling systematic classification and synthesis with improved clarity and unified understanding. Our survey analyzes and compares 29 commonly used datasets and evaluation metrics relevant to these areas.
@@
$\bullet$ Finally, we discuss significant challenges and propose directions for future research, emphasizing the essential contribution of these domains to AI assistant development. Our synthesis aims to inform and inspire future work in the field.
@@ 2 Primer On Egocentric Procedural AI Assistant
This section consists of two parts. First, we divide the construction of the EgoProceAssist into three core areas. Then, we analyze and summarize the relevant techniques.
2.1 Three Core Functions and Tasks
Here, we further delineate the essential capabilities required for EgoProceAssist. First, it must be capable of detecting procedural errors in real time from video streams. This ability extends beyond mere anomaly detection, enabling the assistant to comprehend everyday procedural tasks with potential procedural errors from an egocentric perspective. Second, the assistant should autonomously learn how to perform a procedural task by observing video demonstrations, automatically identifying key steps and their sequences. This capability enables more intelligent and efficient procedural guidance for users. Finally, EgoProceAssist needs to understand egocentric videos and perform relevant question-answering tasks, improving its interactivity and allowing the assistant to provide timely guidance when needed.
After conducting in-depth and comprehensive research on those capabilities, the development of EgoProceAssist can be systematically divided into three core tasks: egocentric procedural error detection, egocentric procedural learning, and egocentric procedural question answering.
@@ 2.2 Related Work
Progressing research in these areas will play a key role in developing EgoProceAssist, as they offer rich opportunities for further innovation. Methods in these domains leverage proven computer vision and deep learning techniques, supporting robust and scalable improvement and deployment.
@@
Figure 2: Overall Structure. In Section II, domains and techniques related to the construction of EgoProceAssist are presented. In Sections III, IV, and V, we provide a comprehensive summary of the existing technical approaches, commonly used datasets, and evaluation metrics for the three core tasks, respectively. For enhanced clarity, we also present comparative tables to highlight performance differences among the methods. Section VI presents experimental investigations assessing the capability of existing models to understand procedural tasks across two distinct domains. Section VII discusses the current challenges faced in the field and explores potential trends for future research and development. Finally, Section VIII offers a comprehensive summary of the findings and conclusions drawn from this work.
@@ 3.1 Definition
Video anomaly detection, which usually operates from a static viewpoint and flags deviations from regular patterns via semantic rules (such as identifying falls), differs fundamentally from procedural error detection. Procedural errors are goal-oriented and require long-term context-sensitive reasoning, rendering standard anomaly detection methods insufficient for identifying such errors.
As shown in Fig. 4, current error detection methods vary in their data usage. Only-video methods rely solely on video, while multimodal approaches utilize multiple data types, such as gaze, to achieve higher accuracy, reduce data requirements, and increase practicality. Both techniques help inform the development of real-time egocentric procedural AI assistants. The performance comparison of the various methods is given in Table 1.
Figure 4: (a) shows the overall framework flowchart summarizing only-video egocentric procedural error detection methods: VQF [patsch2025technical], EgoPED [lee2024error], TI-PREGO [plini2024ti], AMNAR [huang2025modeling]. (b) shows a flowchart summarizing multimodal-based methods: KRR [ding2023every], TAC [storks2024transparent], EgoOops [haneji2024egooops], GC [mazzamuto2025gazing], where some methods use gaze or procedure-related text as additional input.
@@ 3.2 Only-video Methods
Recent advances in procedural error detection include the dual-branch architecture by Flaborea et al. (PREGO) [flaborea2024prego], which takes videos as input, integrates current actions into a symbolic reasoning module, and utilizes large language models (LLMs) for contextual next-action prediction. Building on this, Plini et al. (TI-PREGO) [plini2024ti] introduce Automatic Chain of Thought (ACoT) [zhang2022automatic] and leverage more powerful LLMs, further enhancing predictive accuracy. Huang et al. (AMNAR) [huang2025modeling] propose generating multiple dynamic representations of correct actions, which also take videos as input, enabling robust error detection across varying task executions and environmental contexts. Patsch et al. (VQF) [patsch2025technical] introduce a Video Q-Former-based end-to-end framework for simultaneously detecting sequential and execution errors, providing textual explanations for identified mistakes. Lee et al. (EgoPED) [lee2024error] propose learning multiple prototypes of correct execution and detecting errors by comparing actions against these prototypes.
Specifically, PREGO [flaborea2024prego] introduces a dual-branch architecture for procedural error detection: one branch employs MiniROAD [an2023miniroad] for real-time action detection via RNN, while the other leverages an LLM to predict subsequent actions through symbolic reasoning on historical action sequences, treating actions as symbolic patterns without model fine-tuning. Errors are flagged when branch outputs diverge, requiring training only for the detection branch and allowing the prediction branch to use an off-the-shelf LLM, thereby enhancing recognition of previously unseen errors. Building on this, TI-PREGO [plini2024ti] integrates frame aggregation strategies in the prediction branch, selecting the NOMA approach for stable action sequences, upgrades to Llama-3.1 8B for superior accuracy and reasoning, and incorporates Automatic Chain of Thought (ACoT) [zhang2022automatic] for step-by-step reasoning, achieving improved performance.
AMNAR [huang2025modeling] detects errors via three modules: a pre-trained I3D model [peng2023i3d] for visual feature extraction, an action segmentation model for identifying action boundaries and labels, and PAPB, which uses a task graph and dynamic programming [bellman1966dynamic] to align the executed actions to the reference sequence using the longest common subsequence (LCS) [bergroth2000survey][paterson1994longest], reducing labeling errors and omissions. The current action sequence $s_{t}$ is converted into $s_{t}^{*}$ by the above method. PAPB identifies the valid next actions $C_{t}$ by extracting the child nodes of $s_{t}^{*}$ in $G:$
$C_{t}=(\bigcup_{a\in s_{t}^{*}}A[a])\setminus s_{t}^{*},$ (1)
where $\bigcup_{a\in s^{*}_{t}}A[a]$ aggregates all successors of nodes in $s^{*}_{t}$ from the adjacency list $A$ of task graph $G$, and $\backslash s^{*}_{t}$ excludes already executed actions. In the Representation Reconstruction Block (RRB) module, dilated convolution is applied to the feature $F_{1:ed_{t-1}}=\{f_{i}\}_{i=1}^{ed_{t-1}}$ before time $t$ to obtain $F_{1:ed_{t-1}}^{\mathrm{conv}}$. A local cross-attention mechanism is then employed, where $C_{t}$ is used as the query, and $F_{1:ed_{t-1}}^{\mathrm{conv}}$ serves as both the key and value. For each valid action $y_{t,i}$, we predict a residual $r_{t,i}$ that refines the cluster center $c_{t,i}$ of the corresponding action class. The normal action representation $f_{t,i}^{normal}$ is then computed as:
$r_{t,i}=\mathcal{E}_{\mathrm{res}}(y_{t,i},F_{1:ed_{t-1}}^{\mathrm{conv}}),$ (2)
$f_{t,i}^{normal}=c_{t,i}+r_{t,i},$ (3)
where $\mathcal{E}_{\text{res}}$ denotes the residual prediction operation via cross-attention, and $c_{t,i}$ is the precomputed cluster center for action class $y_{t,i}$, derived from normal training samples. The resulting set of representations, denoted $f_{t}^{\text{normal}}=\{f_{t,i}^{\text{normal}}\}_{i=1}^{|C_{t}|}$, encapsulates all valid next actions:
$f_{t}^{\text{normal}}=\mathcal{E}(C_{t},F_{1:ed_{t-1}}),$ (4)
where $\mathcal{E}$ integrates all operations within this block. Finally, the process proceeds to the Representation Matching Block (RMB), where the Euclidean distance between the current action feature and each normal action representation is computed to determine whether an error has occurred:
$d_{t,i}=\|f_{t}^{\mathrm{action}}-f_{t,i}^{\mathrm{normal}}\|_{2},$ (5)
[…]
VQF [patsch2025technical] is an end-to-end framework utilizing a Vision Transformer (ViT) [dosovitskiy2020image] for per-frame feature extraction and a Video Q-Former [li2023blip], adapted from BLIP-2, to capture spatiotemporal dependencies in video data. The resulting features are classified via a linear layer to predict error logits $m$, with a trigger mechanism activating error explanation only when the sigmoid output exceeds a threshold $\tau$. For the explanation, the spatiotemporally-aware features are linearly projected and input to a large language model (e.g., LLaMA 2 [touvron2023llama]) alongside a prompt, enabling the generation of natural language error explanations.
EgoPED [lee2024error] employs a pre-trained video feature extractor (e.g., I3D [carreira2017quo]) for frame-level feature extraction, and Faster R-CNN to identify object and hand bounding boxes, thus detecting active objects. A 3-layer GCN processes a relation graph to derive relational features, which are input to a CSPL module [lin2022prototypical]. K-means clustering generates step prototypes, and contrastive learning (InfoNCE loss [oord2018representation]) aligns features with corresponding prototypes while distinguishing different steps. The loss function is defined as follows:
$\mathcal{L}_{cspl}=-\sum_{i}\log\left(\frac{s_{n,i}^{+}}{s_{n,i}^{+}+s_{n,i}^{-}}\right),$ (6)
where $s_{n,i}^{+}$ is positive sample similarity and $s_{n,i}^{-}$ is negative sample similarity.
@@
TI-PREGO (Oracle+Llama 3.1) [plini2024ti] arXiv 2024 - - - - 97.8 -
TI-PREGO (MiniRoad+Llama 3.1) [plini2024ti] arXiv 2024 - - - - 97.2 -
@@
TI-PREGO (Oracle+Llama 3.1) [plini2024ti] arXiv 2024 - - - - 100.0 -
TI-PREGO (MiniRoad+Llama 3.1) [plini2024ti] arXiv 2024 - - - - 93.3 -
@@ 3.3 Multimodal-based Methods
Besides only-video approaches, numerous methods integrate multimodal data, such as text for error detection. Some techniques also learn the correct behavioral patterns associated with a given label and compare them with actual executions to identify potential errors.
Ding et al. (KRR) [ding2023every] construct a dynamically updated knowledge base to formalize assembly logic into inferable rules, enabling adaptation to complex scenarios via comparison of current and historical actions. Storks et al. (TAC) [storks2024transparent] utilize Vision Language Models (VLMs) for self-dialogue explanations through a question-answering mechanism, ensuring transparent and interpretable mistake detection. Haneji et al. (EgoOops) [haneji2024egooops] integrate video-text alignment with mistake classification, using procedural text to identify erroneous actions in videos. Mazzamuto et al. (GC) [mazzamuto2025gazing] employ analysis of users’ eye movements to identify attentional deviations as indicators of procedural errors.
Specifically, KRR [ding2023every] employs graph-based spatial and temporal beliefs to assess action feasibility and order. The knowledge base is dynamically updated with observed actions. Error detection utilizes reasoning based on current and stored historical actions via encoded rules. This framework innovatively models assembly error detection as a knowledge representation and reasoning task, yielding strong, interpretable results.
TAC [storks2024transparent] accepts procedural text and a video frame as input, outputting a binary success/error decision accompanied by a self-dialogue rationale, which is an iteratively generated sequence of question-answer pairs reflecting the reasoning process. Users can review these pairs to assess model reliability. The model creates candidate questions based on procedural text and dialogue context, and then determines correctness by synthesizing these questions with the current frame. Explanation rationality is further evaluated by a natural language inference (NLI) model (BART-MNLI) [lewis2019bart], which quantifies the relevance and informativeness of each Q&A pair.
EgoOops [haneji2024egooops] builds on StepFormer [pramanick2023egovlpv2] for self-supervised video-text alignment, introducing StepFormer++ with fully supervised loss and fine-tuning on EgoOops dataset annotations. EgoVLPv2 [cui2019class] replaces UniVL for improved egocentric video feature extraction, followed by temporal pooling and concatenation with text features. A two-layer MLP [kruse2022multi] predicts probabilities for correct, mistake, and correction labels. Error detection occurs offline after task completion, requiring full procedural text and complete video, thus precluding real-time feedback.
GC [mazzamuto2025gazing] utilizes egocentric videos and prior gaze trajectories as inputs, building upon [lai2022eye] by introducing a Transformer-based encoder-decoder architecture with two fusion strategies: channel fusion, which concatenates gaze heatmaps with video frames, and correlation fusion, which applies an attention mechanism to model the interplay between gaze and visual features. The model predicts future gaze trajectories, which are then compared to ground truth to calculate a deviation score. Notably, this method detects procedural errors by analyzing gaze behaviors without relying on error-labeled data, and both fusion strategies contribute to its robustness.
@@
EgoOops [haneji2024egooops] arXiv 2024 6.8h 50 538 5 Mixed Ego

### v2

@@ Abstract
Driven by recent advances in vision-language models (VLMs) and egocentric perception research, the emerging topic of an egocentric procedural AI assistant (EgoProceAssist) is introduced to step-by-step support daily procedural tasks in a first-person view. In this paper, we start by identifying three core tasks in EgoProceAssist: egocentric procedural error detection, egocentric procedural learning, and egocentric procedural question answering, then introduce two enabling dimensions: real-time and streaming video understanding, and proactive interaction in procedural contexts. We define these tasks within a new taxonomy as the EgoProceAssist’s essential functions and illustrate how they can be deployed in real-world scenarios for daily activity assistants. Specifically, our work encompasses a comprehensive review of current techniques, relevant datasets, and evaluation metrics across these five core areas. To clarify the gap between the proposed EgoProceAssist and existing VLM-based assistants, we conduct novel experiments to provide a comprehensive evaluation of representative VLM-based methods. Through these findings and our technical analysis, we discuss the challenges ahead and suggest future research directions. Furthermore, an exhaustive list of this study is publicly available in an active repository that continuously collects the latest work: https://github.com/z1oong/Building-Egocentric-Procedural-AI-Assistant.
@@
Egocentric perception, procedural error detection, procedural learning, procedural question answering, VLM.
@@ 1 Introduction
Procedural tasks, which require adherence to specific sequences of steps, are central to numerous daily and industrial activities. Errors in the order or execution of these steps can lead to inefficiencies, suboptimal outcomes, or even safety risks. Traditional video analysis methods, while capable of visual pattern recognition, lack a deep understanding of video content. They struggle to comprehend logical procedural narratives and face challenges in integrating multifunctional assistance, such as detection and video question answering (VQA). Recent advances in VLM [cheng2024videollama][lin2024video][lillava] and egocentric (first-person view) vision technologies [thatipelli2025egocentric][li2025challenges], combined with augmented and virtual reality devices, have enabled the development of egocentric AI assistants for procedural task guidance [yang2025egolife, huang2025vinci].
@@
Based on the envisioned functions, such as the ability to detect procedural errors in real time, the ability to learn the key steps of procedural tasks independently from videos, the ability to understand egocentric videos and answer relevant questions, this paper identifies three new technical approaches and two enabling dimensions, to address key requirements for building an egocentric procedural AI assistant (EgoProceAssist): egocentric procedural error detection, egocentric procedural learning, egocentric procedural question answering, real-time and streaming video understanding, and proactive interaction in procedural contexts. The main contributions of this work can be summarized as follows:
Figure 1: Taxonomy of EgoProceAssist [flaborea2024prego, plini2025ti, huang2025modeling, patsch2025technical, lee2024error][ding2023every][storks2024transparent, haneji2025egooops, mazzamuto2025gazing, bansal2024united, alayrac2016unsupervised, elhamifar2019unsupervised, kukleva2019unsupervised, chowdhury2024opel, sener2015unsupervised, lin2022learning, zhou2018towards, zhukov2019cross, naing2020procedure, bansal2022my, mahmood2025procedure, shah2023steps, wang2023lifelongmemory, zhang2024hcqa, zhang2025hcqa, taluzzi2025pixels][yang2025egolife][di2024grounded][huang2025vinci][suglia2024alanavlm, chen2025grounded, biswas2025raven, yemmego][10.1145/3610910, 10.1145/3659600, chen2025egoagentjointpredictiveagent, fan2025embodiedvideoagentpersistentmemory, chen2024videollm, qian2025dispider, zhang2024internlm, zhang2024flash, wang2025streambridge, xiongstreaming, qian2024streaming, wang2024videollm, wu2024videollm, yang2025socialmind, bandyopadhyay2025yeti, zhang2024proagent, luproactive, wang2025egosocial, zhang2025proactive, zhao2023antgpt, fang2024egocentric, kulkarni2025egovita, liu2023egocentric, rhinehart2017first, rodin2024action].
@@
$\bullet$ We present novel method taxonomies across the five domains, as shown in Fig. 1, enabling systematic classification and synthesis with improved clarity and a unified understanding. Our survey analyzes 39 commonly used datasets and evaluation metrics relevant to these areas.
@@
$\bullet$ Finally, we discuss significant challenges and propose directions for future research, emphasizing the essential contribution of these domains to AI assistant development. Our synthesis aims to inspire future work in the field.
The remainder of this paper is organized as follows. In Section 2, domains and techniques related to the construction of EgoProceAssist are presented. In Sections 3, 4, 5, and 6, we provide a comprehensive summary of the existing technical approaches, commonly used datasets, and evaluation metrics for the three core tasks and two enabling dimensions, respectively. For enhanced clarity, we also present comparative tables to highlight performance differences among the methods. Section 7 presents experimental investigations assessing the capability of existing models to understand procedural tasks across two distinct domains. Section 8 discusses the current challenges faced in the field and explores potential trends for future research and development. Finally, Section 9 offers a comprehensive summary of the findings and conclusions drawn from this work.
@@ 2 Primer On Egocentric Procedural AI Assistant
This section consists of two parts. We first review the current progress in related fields, then structure EgoProceAssist around three function-oriented domains and two enabling dimensions.
@@ 2.1 Related Work
Progressing research in these areas will play a key role in developing EgoProceAssist, as they offer rich opportunities for further innovation. Methods in these domains leverage proven computer vision and deep learning techniques, enabling robust, scalable improvement and deployment.
AI Assistants on Wearable Devices. As wearable devices become more widely deployed, research on AI assistants built on these platforms has grown. StretchAR [10.1145/3550305] introduces tactile feedback via smart wristbands, and PrISM-QA [10.1145/3699759] uses smartwatches for voice-based conversational assistance. Although these systems effectively exploit wearable devices, they make only limited use of visual input and thus are less suitable for tasks requiring rich real-world perception. LiSee [10.1145/3550282] employs a neck-worn earphone to support visually impaired users in daily object retrieval using combined visual and auditory cues. CASES [10.1145/3610910] runs on AI glasses to assist reading by analyzing visual attention and text. Also on AI glasses, OS-1 [10.1145/3659600] leverages an LLM for conversational tutoring, while EMOShip [Zhao_2022] performs emotion analysis. Vinci [huang2025vinci] integrates visual and language understanding to support daily activities. EgoAgent [chen2025egoagentjointpredictiveagent] predicts future states and human poses from egocentric views, and Embodied VideoAgent [fan2025embodiedvideoagentpersistentmemory] enables memory and interaction in dynamic 3D scenes through multimodal fusion and tool invocation. However, these systems are largely tailored to specific tasks and user groups and lack generalization. Existing AI assistants are evolving from single-modality processing to multimodal understanding, integrating diverse functionalities to enhance cross-task generalization and better support everyday life. In real-world settings, AI-assisted tasks are often procedural: they follow ordered steps, target specific goals, and pose substantial comprehension challenges. Simple QA tools or single-modality LLM reasoning are often insufficient. These limitations highlight the research significance of EgoProceAssist.
@@
2.2 Three Core Functions and Two Enabling Dimensions
In the following sections, we focus on several core functions that EgoProceAssist is designed to fulfill, grounded in potential user concerns and anticipated needs in procedural tasks. Consider, for example, furniture assembly, where textual instructions alone may be insufficient and difficult to follow. EgoProceAssist, capable of providing stepwise guidance, detecting omitted or out‑of‑order actions, and promptly responding to task‑related queries, would be of substantial practical value. This requires the assistant to first conduct procedural error detection, enabling timely identification of errors such as incorrect steps and execution mistakes (e.g., use of inappropriate tools). It must also recognize key action sequences from observed video, thereby endowing the system with autonomous procedural learning. By retrieving demonstration videos of similar tasks, the assistant can learn how to complete the task and provide more effective guidance. Timely feedback on the user’s current step further facilitates goal achievement. Finally, the assistant must support continuous procedural question answering and provide timely guidance. On this basis, we regard Egocentric Procedural Error Detection, Egocentric Procedural Learning, and Egocentric Procedural Question Answering as the three core and indispensable functions of EgoProceAssist.
These three functions are intrinsically interdependent rather than purely parallel. Learning correct action sequences from similar videos and comparing them with observed behavior offers an efficient and accurate mechanism for error detection [huang2025modeling], while simultaneously supplying rich priors for procedural question answering. Conversely, the strong reasoning capabilities of VQA models can improve the reliability of both procedural learning and error detection [flaborea2024prego, plini2025ti]. Moreover, information about incorrect steps obtained via error detection can be used to validate and refine procedural learning and question answering. In subsequent work, we categorize these three fundamental functions according to their functional roles, thereby enabling clearer summarization, classification, and generalization.
While the three core functionalities outlined above define what EgoProceAssist can do, two critical enabling dimensions determine whether the model can be deployed in real-world settings. First, EgoProceAssist must process real-time, streaming video. Only an assistant with low latency and rapid response can provide effective and timely support, but detecting a procedural error five seconds after it occurs offers little practical value. Second, EgoProceAssist must provide proactive guidance and assistance. Existing procedural task-support methods, such as procedural question answering, require users to actively initiate queries. In real-world scenarios where an AI assistant is needed, users often lack the expertise to formulate appropriate questions. Moreover, in urgent error cases, users may be unaware of the problem and thus cannot trigger passive detection or assistance. Consequently, these requirements introduce Real-time and Streaming Video Understanding, and Proactive Interaction in Procedural Contexts as two key enabling dimensions for EgoProceAssist.
@@ 3.1 Definition
Video anomaly detection, which typically operates from a static viewpoint and flags deviations from regular patterns using semantic rules (e.g., identifying falls), differs fundamentally from procedural error detection. Procedural errors are goal-oriented and require long-term context-sensitive reasoning, rendering standard anomaly detection methods insufficient for identifying them.
As shown in Fig. 3, current error detection methods differ in their data usage. Only-video methods rely solely on video, while multimodal approaches utilize multiple data types, such as gaze, to achieve higher accuracy, reduce data requirements, and increase practicality. Both techniques help inform the development of real-time egocentric procedural AI assistants. The performance comparison of the various methods is given in Table 1.
Figure 3: (a) shows the overall framework flowchart summarizing only-video egocentric procedural error detection methods: VQF [patsch2025technical], EgoPED [lee2024error], TI-PREGO [plini2025ti], AMNAR [huang2025modeling]. (b) shows a flowchart summarizing multimodal-based methods: KRR [ding2023every], TAC [storks2024transparent], EgoOops [haneji2025egooops], GC [mazzamuto2025gazing], where some methods use gaze or procedure-related text as additional input.
@@ 3.2 Only-video Methods
Although relying solely on video input, some methods such as Flaborea et al. (PREGO) [flaborea2024prego] and Plini et al. (TI-PREGO) [plini2025ti] leverage the reasoning capabilities of LLMs for contextual next-action prediction, enabling timely responses and the detection of previously unseen errors. The incorporation of Automatic Chain of Thought (ACoT) [zhang2022automatic] reasoning further enhances reasoning ability and improves detection accuracy. However, these approaches are highly dependent on the precision of action recognition and prediction, and may exhibit limited generalization due to the inherent constraints of LLMs. To reduce reliance on prediction accuracy, other methods, such as Huang et al. (AMNAR) [huang2025modeling] and Lee et al. (EgoPED) [lee2024error], generate multiple dynamic representations of correct actions and detect errors by comparing them. Although they cannot provide rapid prediction-based responses or effectively detect unseen errors, they enable robust error detection across diverse task executions and environmental contexts. These methods can accommodate changes in action distributions, handle scenarios with multiple valid actions, and detect various error types, such as omissions and modifications. To improve error interpretability, Patsch et al. (VQF) [patsch2025technical] propose a Video Q-Former–based end-to-end framework that simultaneously detects sequential and execution errors while providing textual explanations for the identified mistakes.
@@
TI-PREGO (Oracle+Llama 3.1) [plini2025ti] CVIU 2025 - - - - 97.8 -
TI-PREGO (MiniRoad+Llama 3.1) [plini2025ti] CVIU 2025 - - - - 97.2 -
@@
TI-PREGO (Oracle+Llama 3.1) [plini2025ti] CVIU 2025 - - - - 100.0 -
TI-PREGO (MiniRoad+Llama 3.1) [plini2025ti] CVIU 2025 - - - - 93.3 -
@@ 3.3 Multimodal-based Methods
Besides only-video approaches, numerous methods integrate multimodal data. Wearable devices nowadays are well-suited for collecting various modal information beyond egocentric visual perception, such as gaze, facial micro-expressions, and audio. Gaze reveals user intentions via attentional patterns, while audio conveys social states and environmental cues that are inaccessible to the egocentric view, and vocal characteristics enable inference of users’ emotions and intentions. However, existing error detection methods only utilize information such as text and gaze. How to better utilize data from other modalities warrants further consideration.
Ding et al. (KRR) [ding2023every] construct a dynamically updated knowledge base that formalizes assembly logic into inferable rules, enabling adaptation to complex scenarios by comparing current and historical actions. Other approaches like Storks et al. (TAC) [storks2024transparent] and Haneji et al. (EgoOops) [haneji2025egooops] integrate textual data to provide transparent and interpretable error detection. However, these methods generally rely on fine-grained dataset annotations, and their performance degrades when generalizing to unseen domains. Meanwhile, video-text-based approaches often suffer from the loss of fine-grained details. Mazzamuto et al. (GC) [mazzamuto2025gazing] addresses these limitations by employing analysis of users’ eye movements. Based on gaze prediction, it does not require access to the full video and can respond rapidly, even to previously unseen error types.
@@
EgoOops [haneji2025egooops] ICCV 2025 6.8h 50 538 5 Mixed Ego